# On the conjugates of Christoffel words

### v3

@@ 9. The Sturmian graph revisited
Moreover, if $i<m-1$ and $k\geq 1$, then for each vertex of the form $UL_{i}^{k}L_{i+1},k\geq 1$, there is an edge labelled $L_{i+1}$ from $U$ to $UL_{i}^{k}L_{i+1}$:
@@
[…]
@@
An example may be useful. Let $m=3,a_{1}=2,a_{2}=1,a_{3}=3$. Then $M_{-1}=b,M_{0}=a,M_{1}=M_{0}^{a_{1}-1}M_{-1}=ab$, $M_{2}=M_{1}^{a_{2}}M_{0}=aba$, $M_{3}=M_{2}^{a_{3}}M_{1}=abaabaabaab$. Since $M_{3}=pab$, we have $p=aba^{2}ba^{2}ba$. The palindromic prefixes of $p$ are $1$, $a$, $aba$, $abaaba$, $p$; the letter following each palindromic prefix is underlined: $p=\underline{a}\,\underline{b}a\underline{a}ba\underline{a}ba$, and therefore $p=Pal(abaa)$. One has $Pal(1)=1,Pal(a)=a,Pal(ab)=aba,Pal(aba)=abaaba$. The tree interpretation is shown in Figure 2: the words $1,a,aba,abaaba,p$ are the nodes on the path from the root to $p$ in the tree of central words. One recovers in two ways that $L_{0}=a,L_{1}=ba,L_{2}=aba$: using $L_{i}=\widetilde{M}_{i}$, or using the last assertion of Proposition 9.4.
$abaabaaba$$abaaba$$aba$$a$$\epsilon$$abaabaaba$$abaaba$$aba$$a$$\epsilon$$a$$ba$$ba$$aba$$aba$$aba$$\frac{4}{7}$$\frac{3}{5}$$\frac{2}{3}$$\frac{1}{2}$$\frac{1}{1}$111122222222333333333333
@@ 10. Appendix: a proof of existence and uniqueness of the greedy and lazy representations
(i) The hypothesis implies $d_{k}=b_{k}$. For $k=0$ and $k=1$, the equality follows from $q_{0}=1$ and $b_{1}=a_{1}-1=q_{1}-1$. Suppose that $k\geq 1$, and that the equality is true for $k-1$ and $k$. Consider an alternating sequence $d_{1},\ldots,d_{k+1}$ with $d_{k+1}\neq 0$; then $d_{k+1}=b_{k+1}$. We have $\sum_{i=1}^{k+1}d_{i}q_{i-1}=b_{k+1}q_{k}+\sum_{i=1}^{k-1}d_{i}q_{i-1}$ (since $d_{k}=0$, because the sequence is alternating) $=a_{k+1}q_{k}+q_{k-1}-1$ (by induction, since $d_{k-1}=b_{k-1}\neq 0$) $=q_{k+1}-1$.
(ii) This is clearly true for $k=0$ and $k=1$, since $q_{0}=1$ and $b_{1}=a_{1}-1=q_{1}-1$. Assume that $k\geq 1$ and that it is true for $k-1$ and $k$, and we prove it for $k+1$; thus we consider a sequence $d_{1},\ldots,d_{k+1}$ with $d_{k+1}=b_{k+1}$. If $d_{k}\neq 0$, then $N=\sum_{i=1}^{k+1}d_{i}q_{i-1}\geq b_{k+1}q_{k}+q_{k-1}=a_{k+1}q_{k}+q_{k-1}=q_{k+1}\geq q_{k+1}-1$. If $d_{k}=0$ then, assuming that $k\geq 2$, we have $d_{k-1}=b_{k-1}$ since the representation is lazy ; thus by induction, $N\geq b_{k+1}q_{k}+q_{k-1}-1=a_{k+1}q_{k}+q_{k-1}-1=q_{k+1}-1$. The remaining case is $k=1,d_{2}=b_{2}=a_{2},d_{1}=0$ and $N=d_{2}q_{1}=a_{2}a_{1}=q_{2}-1$. ∎
@@
(i) Let $0\leq k\leq m$. We prove the existence of a greedy representation $N=\sum_{i=1}^{k}d_{i}q_{i-1}$ for each $N$ satisfying $N\leq q_{k}-1$; by the previous observation, this will prove the existence of a greedy representation for each $N$ with $10\leq N\leq q_{m}-$ 1. For $k=0$, we have $N=0$ and existence is clear. For $k=1$, $N\leq q_{1}-1$; then $N\leq a_{1}-1$ and we have $N=d_{1}q_{0}$, $d_{1}=N\leq a_{1}-1$ and existence follows.
Suppose now that $k\geq 1$, and let $N$ satisfy $N\leq q_{k+1}-1$. Then $N\leq a_{k+1}q_{k}+q_{k-1}-1$. Since $q_{k-1}-1<q_{k}$, we have $N<(a_{k+1}+1)q_{k}$; thus, performing the Euclidean division of $N$ by $q_{k}$, there are uniquely determined $r,t$ with $N=tq_{k}+r$, $10\leq r\leq q_{k}-$ 1 and $t\leq a_{k+1}$.
@@
(ii) We prove now the existence of the lazy representation. We observe that ()(*)( ∗ ) the sequence $q_{k}+q_{k-1}-2$ is strictly increasing for $k=0,\ldots,m$, with first value $-1$. We prove by induction on $k=1,\ldots,m$ that if $N\leq q_{k}+q_{k-1}-2$, then $N$ has a lazy representation of the form $N=\sum_{i=1}^{k}d_{i}q_{i-1}$. For $k=1$, the inequality is $N\leq a_{1}-1$, and we have indeed $N=d_{1}q_{0}$, with $d_{1}=N$, $10\leq d_{1}\leq a_{1}-$ 1. Assume now that $k\geq 1$, and that the property holds for $k$, and we prove it when $N$ satisfies $N\leq q_{k+1}+q_{k}-2$. By induction, we may assume that $q_{k}+q_{k-1}-2<N$. We have $q_{k}+q_{k-1}-1\leq N$ and since $q_{k-1}-1\geq 0$ (because $k\geq 1$), there exists $j$, $1\leq j\leq a_{k+1}$ such that $jq_{k}\leq N$ and we take $j$ maximal. Then either $j=a_{k+1}$ and $N\leq q_{k}+q_{k+1}-2=(j+1)q_{k}+q_{k-1}-2$; or $j<a_{k+1}$ and then $j+1\leq a_{k+1}$ and by maximality, $N<(j+1)q_{k}\leq(j+1)q_{k}+q_{k-1}-1$ and we have $N\leq(j+1)q_{k}+q_{k-1}-2$, too.
Write $N=jq_{k}+N^{\prime}$; then $20\leq N^{\prime}\leq q_{k}+q_{k-1}-$ 2. By induction, $N^{\prime}$ has a lazy representation $N^{\prime}=d_{1}q_{0}+\cdots+d_{k}q_{k-1}.$ Then, $d_{1}q_{0}+\cdots+d_{k}q_{k-1}+jq_{k}$ is a lazy representation of $N$, except when $d_{k}=0$ and $d_{k-1}\not=b_{k-1}$ (so that $k\geq 2$); since $a_{k}=b_{k}$, $d_{1}q_{0}+\cdots+(d_{k-1}+1)q_{k-2}+b_{k}q_{k-1}+(j-1)q_{k}$ is then a lazy representation of $N$.
We prove now uniqueness of the lazy representation. We may assume that $N\neq 0$. Suppose that $N$ has two lazy representations $N=\sum_{i=1}^{k}d_{i}q_{i-1}$, $N=\sum_{i=1}^{h}e_{i}q_{i-1}$, written in such a way that $d_{k}\neq 0\neq e_{h}$. We have by Lemma 10.2 (ii): $q_{k-1}+q_{k-2}-2<N\leq q_{k}+q_{k-1}-2$ and $q_{h-1}+q_{h-2}-2<N\leq q_{h}+q_{h-1}-2$. This implies that $k=h$, by observation ()(*)( ∗ ).

### v4

@@ 9. The Sturmian graph revisited
Moreover, if i<m1i<m-1 and $k\geq 1$, then for each vertex of the form $UL_{i}^{k}L_{i+1},k\geq 1$, there is an edge labelled $L_{i+1}$ from $U$ to $UL_{i}^{k}L_{i+1}$:
@@
There is a simple algorithm to recover the compact graph from the minimal automaton $\mathcal{A}$ of the set of suffixes of $p$: one chooses some vertex $v$ which is not final, which has only one outgoing edge $v\overset{t}{\to}v^{\prime}$; one considers all incoming edges, all labelled by the same letter $z$ (since the automaton is homogeneous); then one suppresses the vertex $v$ and one lets the incoming edges point towards $v^{\prime}$, adding $t$ at the end of their label. Iterating this procedure, called compaction, one recovers $(V,E)$.
@@
An example may be useful. Let $m=3,a_{1}=2,a_{2}=1,a_{3}=3$. Then $M_{-1}=b,M_{0}=a,M_{1}=M_{0}^{a_{1}-1}M_{-1}=ab$, $M_{2}=M_{1}^{a_{2}}M_{0}=aba$, $M_{3}=M_{2}^{a_{3}}M_{1}=abaabaabaab$. Since $M_{3}=pab$, we have $p=aba^{2}ba^{2}ba$. The palindromic prefixes of $p$ are $1$, $a$, abaaba, abaabaabaaba, $p$; the letter following each palindromic prefix is underlined: $p=\underline{a}\,\underline{b}a\underline{a}ba\underline{a}ba$, and therefore $p=Pal(abaa)$. One has Pal(1)=1,Pal(a)=a,Pal(ab)=aba,Pal(aba)=abaabaPal(1)=1,Pal(a)=a,Pal(ab)=aba,Pal(aba)=abaaba. The tree interpretation is shown in Figure 2: the words $1,a,aba,abaaba,p$ are the nodes on the path from the root to $p$ in the tree of central words. One recovers in two ways that $L_{0}=a,L_{1}=ba,L_{2}=aba$: using $L_{i}=\widetilde{M}_{i}$, or using the last assertion of Proposition 9.4.
[…]
@@ 10. Appendix: a proof of existence and uniqueness of the greedy and lazy representations
(i) The hypothesis implies $d_{k}=b_{k}$. For $k=0$ and $k=1$, the equality follows from $q_{0}=1$ and $b_{1}=a_{1}-1=q_{1}-1$. Suppose that $k\geq 1$, and that the equality is true for k1k-1 and $k$. Consider an alternating sequence $d_{1},\ldots,d_{k+1}$ with $d_{k+1}\neq 0$; then $d_{k+1}=b_{k+1}$. We have $\sum_{i=1}^{k+1}d_{i}q_{i-1}=b_{k+1}q_{k}+\sum_{i=1}^{k-1}d_{i}q_{i-1}$ (since $d_{k}=0$, because the sequence is alternating) $a_{k+1}q_{k}+q_{k-1}-1$ (by induction, since $d_{k-1}=b_{k-1}\neq 0$) $q_{k+1}-1$.
(ii) This is clearly true for $k=0$ and $k=1$, since $q_{0}=1$ and $b_{1}=a_{1}-1=q_{1}-1$. Assume that $k\geq 1$ and that it is true for k1k-1 and $k$, and we prove it for $k+1$; thus we consider a sequence $d_{1},\ldots,d_{k+1}$ with $d_{k+1}=b_{k+1}$. If $d_{k}\neq 0$, then $N=\sum_{i=1}^{k+1}d_{i}q_{i-1}\geq b_{k+1}q_{k}+q_{k-1}=a_{k+1}q_{k}+q_{k-1}=q_{k+1}\geq q_{k+1}-1$. If $d_{k}=0$ then, assuming that $k\geq 2$, we have $d_{k-1}=b_{k-1}$ since the representation is lazy ; thus by induction, $N\geq b_{k+1}q_{k}+q_{k-1}-1=a_{k+1}q_{k}+q_{k-1}-1=q_{k+1}-1$. The remaining case is $k=1,d_{2}=b_{2}=a_{2},d_{1}=0$ and $N=d_{2}q_{1}=a_{2}a_{1}=q_{2}-1$. ∎
@@
(i) Let $0\leq k\leq m$. We prove the existence of a greedy representation $N=\sum_{i=1}^{k}d_{i}q_{i-1}$ for each $N$ satisfying $N\leq q_{k}-1$; by the previous observation, this will prove the existence of a greedy representation for each $N$ with $0\leq N\leq q_{m}-1$. For $k=0$, we have $N=0$ and existence is clear. For $k=1$, $N\leq q_{1}-1$; then $N\leq a_{1}-1$ and we have $N=d_{1}q_{0}$, $d_{1}=N\leq a_{1}-1$ and existence follows.
Suppose now that $k\geq 1$, and let $N$ satisfy $N\leq q_{k+1}-1$. Then $N\leq a_{k+1}q_{k}+q_{k-1}-1$. Since $q_{k-1}-1<q_{k}$, we have $N<(a_{k+1}+1)q_{k}$; thus, performing the Euclidean division of $N$ by $q_{k}$, there are uniquely determined $r,t$ with $N=tq_{k}+r$, $0\leq r\leq q_{k}-1$ and $t\leq a_{k+1}$.
@@
(ii) We prove now the existence of the lazy representation. We observe that ()(*) the sequence $q_{k}+q_{k-1}-2$ is strictly increasing for $k=0,\ldots,m$, with first value 1-1. We prove by induction on $k=1,\ldots,m$ that if $N\leq q_{k}+q_{k-1}-2$, then $N$ has a lazy representation of the form $N=\sum_{i=1}^{k}d_{i}q_{i-1}$. For $k=1$, the inequality is $N\leq a_{1}-1$, and we have indeed $N=d_{1}q_{0}$, with $d_{1}=N$, $0\leq d_{1}\leq a_{1}-1$. Assume now that $k\geq 1$, and that the property holds for $k$, and we prove it when $N$ satisfies $N\leq q_{k+1}+q_{k}-2$. By induction, we may assume that $q_{k}+q_{k-1}-2<N$. We have $q_{k}+q_{k-1}-1\leq N$ and since $q_{k-1}-1\geq 0$ (because $k\geq 1$), there exists $j$, $1\leq j\leq a_{k+1}$ such that $jq_{k}\leq N$ and we take $j$ maximal. Then either $j=a_{k+1}$ and $N\leq q_{k}+q_{k+1}-2=(j+1)q_{k}+q_{k-1}-2$; or $j<a_{k+1}$ and then $j+1\leq a_{k+1}$ and by maximality, $N<(j+1)q_{k}\leq(j+1)q_{k}+q_{k-1}-1$ and we have $N\leq(j+1)q_{k}+q_{k-1}-2$, too.
Write $N=jq_{k}+N^{\prime}$; then $0\leq N^{\prime}\leq q_{k}+q_{k-1}-2$. By induction, $N^{\prime}$ has a lazy representation $N^{\prime}=d_{1}q_{0}+\cdots+d_{k}q_{k-1}.$ Then, $d_{1}q_{0}+\cdots+d_{k}q_{k-1}+jq_{k}$ is a lazy representation of $N$, except when $d_{k}=0$ and $d_{k-1}\not=b_{k-1}$ (so that $k\geq 2$); since $a_{k}=b_{k}$, $d_{1}q_{0}+\cdots+(d_{k-1}+1)q_{k-2}+b_{k}q_{k-1}+(j-1)q_{k}$ is then a lazy representation of $N$.
We prove now uniqueness of the lazy representation. We may assume that $N\neq 0$. Suppose that $N$ has two lazy representations $N=\sum_{i=1}^{k}d_{i}q_{i-1}$, $N=\sum_{i=1}^{h}e_{i}q_{i-1}$, written in such a way that $d_{k}\neq 0\neq e_{h}$. We have by Lemma 10.2 (ii): $q_{k-1}+q_{k-2}-2<N\leq q_{k}+q_{k-1}-2$ and $q_{h-1}+q_{h-2}-2<N\leq q_{h}+q_{h-1}-2$. This implies that $k=h$, by observation ()(*).